OpenStudy (anonymous):

Is it possible to rewrite x(-1)^x-(-1)^x=y(-1)^y-(-1)^y In such a way i could reduce it to x=y ?

OpenStudy (anonymous):

\[x,y \ge0\]

myininaya (myininaya):

is that? \[x(-1)^x-(-1)^x=y(-1)^y-(-1)^y\] ?

hero (hero):

I'd factor out \((-1)^x\) and \((-1)^y\) first

hero (hero):

Because if you did that first, then you'd have \((-1)^x(x - 1) = (-1)^y(y - 1)\)

hero (hero):

Afterwards, you can apply all the stuff @myininaya was talking about

hero (hero):

After applying all those steps, you should end with the desired result.

OpenStudy (anonymous):

thats where i started from

hero (hero):

Can you show all the steps you have completed up to this point?

OpenStudy (anonymous):

\[(-1)^x(x-1)=(-1)^y(y-1)\] that is

OpenStudy (anonymous):

sure

hero (hero):

Apparently only two of the four cases presented by @myininaya actually work.

OpenStudy (anonymous):

Proving injection by contrapositive method \[f(x)=f(y)\] \[\frac{ ((-1)^x(2x-1)+1) }{ 4 }=\frac{ ((-1)^y(2y-1)+1) }{ 4 }\] . . . . \[(-1)^x(x-1)=(-1)(y-1)\] is as far as i've come. I don't know if there's another way to prove it with sets f: N -> Z

OpenStudy (anonymous):

multiplied both sides with 4 subtracted 1 on both sides expaneded the brackets divided each expression with 2

OpenStudy (anonymous):

ect...

OpenStudy (anonymous):

etc*

hero (hero):

How did you conclude that \( ((-1)^x(2x-1)+1)=((-1)^y(2y-1)+1)\) is equivalent to \((-1)^x(x-1)=(-1)(y-1)\) What did you do with the 2?

OpenStudy (anonymous):

\[2x(-1)^x-(-1)^x=2y(-1)^y-(-1)^y\] \[<->\] \[\frac{ 2y(-1)^x-(-1)^x }{ 2 }=\frac{ 2y(-1)^y-(-1)^y }{ 2 }\]

OpenStudy (anonymous):

ah i see hehe

OpenStudy (anonymous):

\[x(-1)^x-(-1)^x=y(-1)^y-(-1)^y\] <-> \[(-1)^x(x-1)=(-1)^y(y-1)\]

OpenStudy (anonymous):

can't see where i should go from here : / (-1)^y,x is annoying me

hero (hero):

Based on what you posted originally, you could simplify it to \((-1)^x(2x - 1) = (-1)^y(2x - 1)\) I don't know how you got it to the other form you have. What you have is not an eqivalent form of the original equation.

OpenStudy (anonymous):

Okay, i stil don't have an idea on what to do, from that expression. \[(-1)^x(2x-1)=(-1)^y(2y-1)\]

OpenStudy (anonymous):

atm i can't see how i should simplify it, without disturbing either one of the sides

myininaya (myininaya):

Well I deleted my cases because I didn't know if x and y were actually integers.

OpenStudy (anonymous):

ye, the original function maps from N to Z

myininaya (myininaya):

go with applying the cases like hero said 2 of the 4 will work those cases being both odd or both even

myininaya (myininaya):

of course you won't have the function is invective if you assume one is even and one is odd because that would imply you don't have f(a)=f(b) implies a=b because there is no way an even can equal an odd

myininaya (myininaya):

injective*

OpenStudy (anonymous):

I see, i think i have to prove it in another way

OpenStudy (anonymous):

but thank you very much :)!